# Counting points on generic character varieties

### v2

@@ 1.1. Overview
This direct relationship between additive and multiplicative character varieties is lost in higher genus; nevertheless, it was conjectured in [HLRV11] that the cohomology of the additive character variety captures the pure part of the cohomology of the multiplicative character variety in all genus. As evidence, the authors proved that additive character varieties are pure. This result relied on the beautiful bridge built by Crawley-Boevey between additive character varieties and quiver varieties associated to comet-shaped quivers [CBIndecomposable]. Under this correspondence, the number of points on the additive character variety matches the Kac polynomials of the corresponding quiver. Utilising these ideas, Hausel, Letellier, and Rodriguez-Villegas proved Kac’s conjecture for arbitrary quivers [HLVKacAnnals2013].
@@
In this paper, we count points on both multiplicative and additive character varieties associated with punctured surface groups and regular semisimple conjugacy classes in arbitrary reductive groups. Our work builds upon and generalises the contributions of Cambò [Cambo17], who studied $\mathrm{Sp}_{2n}$-character varieties; [BK22], who investigated character stacks associated with compact surface groups; and [KNP23], who examined character varieties in presence of regular unipotent conjugacy classes at the punctures.
One of our main results is that the number of points of multiplicative character varieties over finite fields are palindromic polynomials. This hints at a $P=W$ conjecture, a curious Poincaré duality, and a curious Hard Lefschetz property for these varieties. [^1] To make further progress in this direction, one needs a reductive analogue of Markman’s result [MarkmanGenerators2001] that the tautological classes (along with those coming from the underlying surface) generate the cohomology. Another important contribution of this paper is an explicit polynomial formula for the number of points of additive character varieties. The resulting polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers. Computer calculations suggest that these polynomials have non-negative coefficients [^2] See [Bailey] for the computer code written using CHEVIE [GHLMP96].; however, lack of direct connection with quiver varieties makes proving this a nontrivial task.
Finally, one of the main contributions of [HRV08, HLRV11] was to relate the number of points of $\mathrm{GL}_{n}$-character varieties with Macdonald polynomials and use this to provide a conjectural formula for the mixed Hodge polynomials of these varieties. [^3] As far as we know, these conjectures (as well as the purity conjecture mentioned above) are still open. Formulating analogous conjectures in the reductive setting is an interesting open problem.
@@ 1.2. Generic character varieties
As explained in §3.2, the generic assumption ensures that the action of $G/Z$ on $R$ has finite étale stabilisers and that $R$ is smooth and equidimensional. Thus, $[X]$ is a smooth Deligne–Mumford stack. Moreover, $X$ and $[X]$ are equidimensional of pure dimension
@@
Let $\overline{X}$ denote the base change of $X$ to $\overline{\mathbb{F}_{q}}$ and $E(\overline{X})$ the $E$-polynomial of $\overline{X}$. The previous theorem implies that $E(\overline{X})=f_{X}$. By analysing the polynomial $f_{X}$, we will obtain:
@@
    If $g>1$, or $g=1$ and $Z$ is nontrivial, then $\chi(\overline{X})=0$.
@@
The analogue of ${X}$ over the complex numbers is Hodge–Tate, and satisfies the curious Poincaré duality and curious hard Lefschetz. Moreover, there is a version of the $P=W$ conjecture for ${X}$.
When $G=\mathrm{GL}_{n}$, Mellit constructed a cell decomposition for ${X}$, thus providing a conceptual explanation of the fact that ${X}$ is polynomial count and proving in addition that it is Hodge–Tate and satisfies curious Hard Lefschetz and curious Poincaré duality [MellitDecomposition]. We expect that ${X}$ has an analogous cell decomposition for arbitrary reductive group $G$.
@@
    The counting polynomial $f_{X}$ depends only on $(G,g,n)$; i.e., it is independent of the conjugacy classes $C_{i}$ (provided they satisfy Assumption 1).
@@
    There are two reasons for the finite base change in the main theorem. First, the inclusion $[G(\mathbb{F}_{q}),G(\mathbb{F}_{q})]\subseteq[G,G](\mathbb{F}_{q})$ may be proper (e.g. for $G=\mathrm{PGL}_{n}$). However, given $S\in[G,G](\mathbb{F}_{q})$, there exists a positive integer $m$, such that $S\in[G(\mathbb{F}_{q^{m}}),G(\mathbb{F}_{q^{m}})]$. Second, if $W^{\prime}$ is a reflection subgroup of $W$, then $T^{W^{\prime}}$ may be disconnected. The group of components $\pi_{0}=\pi_{0}(T^{W^{\prime}})$ is a finite étale group which may be non-split (i.e. the action of Galois group may be nontrivial). This implies that $|\pi_{0}(\mathbb{F}_{q})|$ depends on the residue of $q$ modulo $|\pi_{0}(\overline{\mathbb{F}_{q}})|$. However, as we shall see, the fact that $\mathrm{char}(k)$ is good for $G$ implies that $q$ is co-prime to $|\pi_{0}(\overline{\mathbb{F}_{q}})|$. Thus, there exists a positive integer $n$ such that
@@ 1.3. The additive analogues
The group $G$ acts on ${A}$ by the adjoint action. We call the resulting GIT quotient
${Y}:={A}/\!\!/G$
@@
We will be working under the obvious additive analogues of Assumption 1. Then the action of $G/Z$ on ${A}$ has finite étale stabilisers, ${A}$ is smooth and equidimensional, and ${Y}$ is also equidimensional of the same dimension as ${X}$.
@@
Let $\overline{Y}$ be the base change of $Y$ to $\overline{\mathbb{F}_{q}}$. Analysing the counting polynomial $f_{Y}$, we find:
@@
    If $G$ is simple of rank less than $5$, $g\leq 10$ and $1\leq n\leq 1000$, then the coefficients of the counting polynomial $f_{Y}$ are non-negative.
Part (b) of this corollary is established with code written using CHEVIE [GHLMP96]. See [Bailey] for the details of this code. The counting polynomial can also be used to give an expression for the Euler characteristic, see §7.4.
@@ 2. Recollections on root systems and reductive groups
In this section, we recall basic facts and notations about root systems and reductive groups. The material here is standard, though perhaps the facts about pseudo-Levi subgroups and quasi-isolated semisimple elements are not as widely known. For these notions, we refer the reader to e.g. [MS03], [Bonnafe05], or [Taylor22].
@@ 2.1. Root systems
[…]
@@ 2.1.6. Isolated pseudo-Levis
$B_{n}$ $(n\geq 3)$ $B_{n}$,  $D_{n}$,  $D_{n-1}\times A_{1}^{\prime}$,  $D_{r}\times B_{n-r}$ ($22\leq r\leq n-$ 2)
@@ 2.2.3. Parabolic subgroups
    If $l\in L$, then $l\in[L,L]$ if and only if $l\in[P,P]$. This follows from the fact that $[P,P]=[L,L]\ltimes N$.
@@ 2.2.8. Pretty good primes
    Suppose $\tilde{G}$ is a possibly disconnected reductive group with connected component $G$, the prime $p=\mathrm{char}(k)$ is pretty good for $G$, and $\pi_{0}(\widetilde{G})$ is étale. Then the centraliser $C_{\tilde{G}}(H)$ of every closed subgroup scheme $H\subseteq\tilde{G}$ is smooth.
@@ 3.1.6. Existence
The above lemma shows that generic conjugacy classes exist. Indeed, if $L$ is a proper standard levi subgroup, then $T\cap[L,L]$ is a proper closed subvariety of $T$. (At the level of Lie algebras, $\mathfrak{t}=\mathfrak{t}\cap[\mathfrak{l},\mathfrak{l}]\oplus Z(\mathfrak{l})$ for $\mathfrak{t}=\mathrm{Lie}(T)$ and the centre $Z(\mathfrak{l})$ is nontrivial if $\mathfrak{l}$ is a proper Levi subalgebra of $\mathfrak{g}$.) Since there are only finitely many standard Levis $L$ and finitely many $w$’s, we conclude that the generic locus of $T^{n}$ is open dense.
@@ 3.2. Generic character varieties
In this section, we assume that $p=\mathrm{char}(k)$ is a very good prime for $G$ and $Z=Z(G)$ is connected. Let $C=(C_{1},\ldots,C_{n})$ be a generic tuple of conjugacy classes of $G$ satisfying $\prod_{i=1}^{n}C_{i}\subseteq[G,G]$. Let $M:G^{2g}\times C_{1}\times\ldots\times C_{n}\rightarrow[G,G]$ be the map
@@
Let ${r=(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in R}$ and suppose $P$ is a parabolic subgroup of $G$ containing all $a_{i},b_{i},$ and $c_{i}$’s. Then, $[P,P]$ contains $[a_{1},b_{1}]\cdots[a_{g},b_{g}]$; thus, $[P,P]$ also contains $\prod_{i=1}^{n}c_{i}$. As $C$ is generic, it follows that $P=G$. Thus, $r$ is irreducible and Schur’s lemma (cf. [Sikora, §4]) implies that $\mathrm{Stab}_{G/Z}(r)$ is finite. Note that this stabiliser is the same as the centraliser $C_{G}(H)$, where $H:=\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle\subseteq G$. By Theorem 10, this group is smooth. Since we already have shown that it is finite, it follows that it is finite and étale.
Now stability follows from the fact that whenever a connected reductive group acts on an affine scheme of finite type with finite stabilisers, then every point is stable, cf. [Mumford, Proposition 0.8]. (Alternatively, one can use the Hilbert–Mumford criterion as in [Boalch14].) The statements about $R$ follow from the fact that $1$ is a regular value for $M$, see [KNP23, §2] for details. ∎
@@
Next, let ${X}:=R/\!\!/G$ and $[{X}]:=[R/(G/Z)]$ denote, respectively, the (generic) character variety and character stack associated to $(\Gamma_{g,n},G,C)$. The previous theorem implies:
Corollary 16.
$[{X}]$ is a smooth Deligne-Mumford stack with coarse moduli space ${X}$.
@@ 3.3. Frobenius Mass Formula
We now work over $k=\overline{\mathbb{F}_{q}}$ and assume that $G$ is equipped with an untwisted Frobenius automorphism $F:G\rightarrow G$. We also assume that the conjugacy classes $C=(C_{1},\ldots,C_{n})$ are $F$-stable. Then the action of the Frobenius $F$ on $G$ and $C_{i}$’s, gives an action of $F$ on $R$, ${X}$ and $[{X}]$. We can therefore speak about the Frobenius fixed points $R^{F}$, ${X}^{F}$, and $[{X}]^{F}$. The first two are finite sets while the last one is a finite groupoid. By definition,
$R^{F}=\Big{\{}(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in(G^{F})^{2g}\times\prod_{i=1}^{n}C_{i}^{F}\,|\,M(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})=1\Big{\}}.$
@@ 3.3.4. Points of character variety
We now assume that the tuple $(C_{1},\ldots,C_{n})$ is generic. We saw above that ${X}$ is a coarse moduli space for the smooth Deligne-Mumford stack $[{X}]$. This implies that $[{X}]$ and $X$ have the same number of points over finite fields, cf. [Behrend91, §2] or [BogaartEdixhoven]. Thus, we conclude:
@@ 3.4. Additive character varieties
The above considerations have additive analogues which we leave to the reader to formulate. The only nontrivial modification is the additive analogue of the Frobenius Mass Formula, which we now discuss.
@@
Let $\mathcal{O}_{1},\ldots,\mathcal{O}_{n}$ be $F$-stable adjoint $G$-orbits in $\mathfrak{g}$. Let $A$ denote the corresponding additive representation variety. Then
${A}^{F}=\Big{\{}(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in(\mathfrak{g}^{F})^{2g}\times\prod_{i=1}^{n}\mathcal{O}_{i}^{F}\,|\,m(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})=0\Big{\}},$
where $m:\mathfrak{g}^{2g}\times\mathcal{O}_{1}\times\ldots\times\mathcal{O}_{n}\rightarrow[\mathfrak{g},\mathfrak{g}]$ is defined by
$m(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})=\sum_{i=1}^{g}[a_{i},b_{i}]+\sum_{i=1}^{n}c_{i}.$
@@
If $\mathcal{O}_{i}$ are generic and stable, then

### v3

@@ 1.1. Overview
This direct relationship between additive and multiplicative character varieties is lost in higher genus; nevertheless, it was conjectured in [HLRV11] that the cohomology of the additive character variety captures the pure part of the cohomology of the multiplicative character variety in all genus. As evidence, the authors proved that additive character varieties are pure, i.e., their cohomology carries a pure Hodge structure. This result relied on the beautiful bridge built by Crawley-Boevey between additive character varieties and quiver varieties associated to comet-shaped quivers [CBIndecomposable]. Under this correspondence, the number of points on the additive character variety matches the Kac polynomials of the corresponding quiver. Utilising these ideas, Hausel, Letellier, and Rodriguez-Villegas proved Kac’s conjecture for arbitrary quivers [HLVKacAnnals2013].
@@
In this paper, we count points on both multiplicative and additive character varieties associated with punctured surface groups and regular semisimple conjugacy classes in arbitrary reductive groups. Our work builds upon and generalises the contributions of Cambò [Cambo17], who studied $\mathrm{Sp}_{2n}$-character varieties; [BK22], who investigated character stacks associated with compact surface groups; and [KNP25], who examined character varieties in presence of regular unipotent conjugacy classes at the punctures.
One of our main results is that the number of points of multiplicative character varieties over finite fields are palindromic polynomials. This hints at a $P=W$ conjecture, a curious Poincaré duality, and a curious Hard Lefschetz property for these varieties. [^1] To make further progress in this direction, one needs a reductive analogue of Markman’s result [MarkmanGenerators2001] that the tautological classes (and those coming from the underlying surface) generate the cohomology. Another important contribution of this paper is an explicit polynomial formula for the number of points of additive character varieties. The resulting polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers. Calculations suggest these polynomials have non-negative coefficients [^2] See [Bailey] for the computer code written using CHEVIE [GHLMP96].; however, lack of direct connection with quiver varieties makes proving this a non-trivial task.
Finally, one of the main contributions of [HRV08, HLRV11] was to relate the number of points of $\mathrm{GL}_{n}$-character varieties with Macdonald polynomials and use this to provide a conjectural formula for the mixed Hodge polynomials of these varieties. [^3] As far as we know, these conjectures (and the purity conjecture mentioned above) are still open. Formulating analogous conjectures in the reductive setting is an interesting open problem.
@@ 1.2. Generic character varieties
As explained in §3.2, the generic assumption ensures that the action of $G/Z$ on $R$ has finite étale stabilisers and that $R$ is smooth and equidimensional. Thus, $[X]$ is a smooth Deligne–Mumford stack. Moreover, $X$ and $[X]$ have the same number of points over a finite field, and are equidimensional of pure dimension
@@
Let $\overline{X}$ denote the base change of $X$ to $\overline{\mathbb{F}_{q}}$ and $E(\overline{X})$ the $E$-polynomial of $\overline{X}$. The previous theorem implies that $E(\overline{X})=f_{X}$. By analysing the counting polynomial $f_{X}$, we will obtain:
@@
    If $g>1$, or $g=1$ and $Z$ is non-trivial, then $\chi(\overline{X})=0$.
@@
The analogue of X{X} over the complex numbers is Hodge–Tate, and satisfies the curious Poincaré duality and curious hard Lefschetz. Moreover, there is a version of the $P=W$ conjecture for X{X}.
When $G=\mathrm{GL}_{n}$, Mellit constructed a cell decomposition for X{X}, thus providing a conceptual explanation of the fact that X{X} is polynomial count and proving in addition that it is Hodge–Tate and satisfies curious Hard Lefschetz and curious Poincaré duality [MellitDecomposition]. We expect that X{X} has an analogous cell decomposition for arbitrary reductive group $G$.
@@
    The polynomial $f_{X}$ depends only on $(G,g,n)$; i.e., it is independent of the conjugacy classes $C_{i}$ (provided they satisfy Assumption 1).
@@
    There are two reasons for the finite base change in the main theorem. First, the inclusion $G(\mathbb{F}_{q}),G(\mathbb{F}_{q})]\subseteq[G,G](\mathbb{F}_{q})$ may be proper (e.g. for $G=\mathrm{PGL}_{n}$). However, given $S\in[G,G](\mathbb{F}_{q})$, there exists a positive integer $m$, such that $S\in[G(\mathbb{F}_{q^{m}}),G(\mathbb{F}_{q^{m}})]$. Second, if $W^{\prime}$ is a reflection subgroup of $W$, then $T^{W^{\prime}}$ may be disconnected. The group of components $\pi_{0}=\pi_{0}(T^{W^{\prime}})$ is a finite étale group which may be non-split (i.e. the action of Galois group may be non-trivial). This implies that $|\pi_{0}(\mathbb{F}_{q})|$ depends on the residue of $q$ modulo $|\pi_{0}(\overline{\mathbb{F}_{q}})|$. However, as we shall see, the fact that $\mathrm{char}(k)$ is good for $G$ implies that $q$ is co-prime to $|\pi_{0}(\overline{\mathbb{F}_{q}})|$. Thus, there exists a positive integer $n$ such that
@@ 1.3. The additive analogues
The group $G$ acts on A{A} by the adjoint action. We call the resulting GIT quotient
Y:=A//G{Y}:={A}/\!\!/G
@@
We will be working under the obvious additive analogues of Assumption 1. Then the action of $G/Z$ on A{A} has finite étale stabilisers, A{A} is smooth and equidimensional, and Y{Y} is also equidimensional of the same dimension as X{X}.
@@
Let $\overline{Y}$ be the base change of $Y$ to $\overline{\mathbb{F}_{q}}$. By analysing the counting polynomial $f_{Y}$, we find:
@@
    If $G$ is simple of rank less than $7$, $0\leq g\leq 10$ and $1\leq n\leq 1000$, then the coefficients of the counting polynomial $f_{Y}$ are non-negative.
Part (b) of this corollary is established with code written using CHEVIE [GHLMP96]; see [Bailey] for the relevant code. The counting polynomial can also be used to give expressions for the Euler characteristic; see §7.4.
@@ 2. Recollections on root systems and reductive groups
In this section, we recall basic facts and notations about root systems and reductive groups. The material here is standard, though perhaps the facts about pseudo-Levi subgroups and quasi-isolated semisimple elements are not as widely known. For these notions, we refer the reader to [MS03], [Bonnafe05], and [Taylor22].
@@ 2.1. Root systems
Let $\Phi$ be a (reduced crystallographic) root system in a finite dimensional Euclidean vector space (V,(.,.))(V,(.\ ,.)). We assume that $\Phi$ spans $V$ and fix a base $\Delta$ of $\Phi$. Then $\Delta$ is a basis of the vector space $V$. Let $W$ be the Weyl group generated by the simple reflections $s_{\alpha}$, $\alpha\in\Delta$.
@@ 2.1.6. Isolated pseudo-Levis
$B_{n}$ $(n\geq 3)$ $B_{n}$, $D_{n}$, $D_{n-1}\times A_{1}^{\prime}$, $D_{r}\times B_{n-r}$ ($2\leq r\leq n-2$)
@@ 2.2.3. Parabolic subgroups
    If $l\in L$, then $l\in[L,L]$ if and only if $l\in[P,P]$. This follows from the fact that $P,P]=[L,L]\ltimes N$.
@@ 2.2.8. Pretty good primes
[…]
@@ 3.1.6. Existence
The above lemma shows that generic conjugacy classes exist. Indeed, if $L$ is a proper standard levi subgroup, then $T\cap[L,L]$ is a proper closed subvariety of $T$. (At the level of Lie algebras, $\mathfrak{t}=\mathfrak{t}\cap[\mathfrak{l},\mathfrak{l}]\oplus Z(\mathfrak{l})$ for $\mathfrak{t}=\mathrm{Lie}(T)$ and the centre $Z(\mathfrak{l})$ is non-trivial if $\mathfrak{l}$ is a proper Levi subalgebra of $\mathfrak{g}$.) Since there are only finitely many standard Levis $L$ and finitely many $w$’s, we conclude that the generic locus of $T^{n}$ is open dense.
@@ 3.2. Generic character varieties
In this section, we assume that $p=\mathrm{char}(k)$ is a very good prime for $G$ and $Z=Z(G)$ is connected. Let $C=(C_{1},\ldots,C_{n})$ be a generic tuple of conjugacy classes of $G$ satisfying $\prod_{i=1}^{n}C_{i}\subseteq[G,G]$. For ease of notation, let $M\colon G^{2g}\times C_{1}\times\cdots\times C_{n}\rightarrow[G,G]$ be the map
@@
Let $r=(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in R$} and suppose $P$ is a parabolic subgroup of $G$ containing all $a_{i},b_{i},$ and $c_{i}$’s. Then, $[P,P]$ contains $a_{1},b_{1}]\cdots[a_{g},b_{g}]$; thus, $[P,P]$ also contains $\prod_{i=1}^{n}c_{i}$. As $C$ is generic, it follows that $P=G$. Thus, $r$ is irreducible and Schur’s lemma (cf. [Sikora, §4]) implies that $\mathrm{Stab}_{G/Z}(r)$ is finite. Note that this stabiliser is the same as the centraliser $C_{G}(H)$, where $H:=\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle\subseteq G$. By Theorem 10, this group is smooth. Since we already have shown that it is finite, it follows that it is finite and étale.
Now stability follows from the fact that whenever a connected reductive group acts on an affine scheme of finite type with finite stabilisers, then every point is stable, cf. [Mumford, Proposition 0.8]. (Alternatively, one can use the Hilbert–Mumford criterion as in [Boalch14].) The statements about $R$ follow from the fact that $1$ is a regular value for $M$, see [KNP25, §2] for details. ∎
@@
Next, let X:=R//G{X}:=R/\!\!/G and [X]:=[R/(G/Z)][{X}]:=[R/(G/Z)] denote, respectively, the (generic) character variety and character stack associated to $(\Gamma_{g,n},G,C)$. The previous theorem implies:
Proposition 16.
[X][{X}] is a smooth Deligne-Mumford stack with coarse moduli space X{X}.
@@ 3.3. Frobenius Mass Formula
We now work over $k=\overline{\mathbb{F}_{q}}$ and assume that $G$ is equipped with an untwisted Frobenius automorphism $F:G\rightarrow G$. We also assume that the conjugacy classes $C=(C_{1},\ldots,C_{n})$ are $F$-stable. Then the action of the Frobenius $F$ on $G$ and $C_{i}$’s, gives an action of $F$ on $R$, X{X} and [X][{X}]. We can therefore speak about the Frobenius fixed points $R^{F}$, ${X}^{F}$, and ${X}]^{F}$. The first two are finite sets while the last one is a finite groupoid. By definition,
$R^{F}=\Big\{(a_{1},\ldots,b_{g},c_{1},\ldots,c_{n})\in(G^{F})^{2g}\times\prod_{i=1}^{n}C_{i}^{F}\,\Big|\,M(a_{1},\ldots,b_{g},c_{1},\ldots,c_{n})=1\Big\}.$
@@ 3.3.4. Points of character variety
We now assume that the tuple $(C_{1},\ldots,C_{n})$ is generic. We saw above that X{X} is a coarse moduli space for the smooth Deligne-Mumford stack [X][{X}]. This implies that [X][{X}] and $X$ have the same number of points over finite fields, cf. [Behrend91, §2] or [BogaartEdixhoven]. Thus, we conclude:
@@ 3.4. Additive character varieties
The above considerations have additive analogues which we leave to the reader to formulate. The only non-trivial modification is the additive analogue of the Frobenius Mass Formula, which we now discuss.
@@
Let $\mathcal{O}_{1},\ldots,\mathcal{O}_{n}$ be $F$-stable adjoint $G$-orbits in $\mathfrak{g}$. Let $A$ denote the corresponding additive representation variety. For ease of notation, let $m\colon\mathfrak{g}^{2g}\times\mathcal{O}_{1}\times\cdots\times\mathcal{O}_{n}\rightarrow[\mathfrak{g},\mathfrak{g}]$ be the map
$m(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}):=\sum_{i=1}^{g}[a_{i},b_{i}]+\sum_{i=1}^{n}c_{i}.$
Then, we have
${A}^{F}=\Big\{(a_{1},\ldots,b_{g},c_{1},\ldots,c_{n})\in(\mathfrak{g}^{F})^{2g}\times\prod_{i=1}^{n}\mathcal{O}_{i}^{F}\,\Big|\,m(a_{1},\ldots,b_{g},c_{1},\ldots,c_{n})=0\Big\}.$
@@
If $\mathcal{O}_{i}$’s are generic and stable, then